### v1

@@ 5.2 Experimentally relevant search on the Y-maze arena
We now consider a model which faithfully represents the domain size and boundaries; food and nest locations; and the number of ants in the system for each experimental realisation. Since the ant population in the arena was dynamic, with ants constantly entering and leaving the nest, to estimate the empirical number of ants for each experiment, we consider the number of ants exploring the arena at the time when any food target was found (end of exploration phase) as the static number $\mathcal{N}$ for the model. The experimental set-up is modelled as a honeycomb lattice confined in a rectangular space, with hard, ‘bouncing’ reflecting walls, which causes persistent walkers to change direction at the boundary and continue its motion in the opposite direction marris2024persistent .
@@
To implement this zig-zag reflecting boundary condition via a fully analytical procedure is a formidable task when the movement is not diffusive. To proceed we use a semi-analytical iterative procedure marris2024persistent ; gueneau2024siegmund , which has been used to study persistent motion with complicated boundary conditions where the analytical representation of the occupation probability is not available. This procedure requires the decomposition of the movement dynamics into matrices that govern only the jumps from each state $i$ to every other state $j$, denoted by $\mathbb{M}(i,j)$. In our case this leads to 36 sparse transition matrices, of size $N_{1}N_{2}\times N_{1}N_{2}$. More concretely, the matrices that govern movement in the bulk have $1/N_{1}N_{2}$ ($\approx 0.31\%$) non-zero elements, while those governing movement on the boundary have $1/N_{1}^{2}N_{2}$ ($\approx 0.022\%$) or $1/N_{1}N_{2}^{2}$ ($\approx 0.0135\%$) non-zero elements. Hence, all matrices benefit from the use of sparse methods leading to efficient computation. See Appendix B for a detailed discussion of the transition matrices.
Encoded into $\mathbb{M}(i,j)$ is also the probability of absorption, i.e. the outgoing probabilities of the trap site, i.e. the site in which the first food source is found, sum up to $1-\rho$, where $\rho$ is the probability of absorption (finding the food) at the target. Experimentally, we estimated $\rho$ by measuring the result of each ant-food interaction. Namely, given an interaction event (i.e. an ant finds a food item), we consider a success (absorption) the ant collecting the food item, or a failure otherwise. Subsequent interactions were considered independent when the ant had walked away from the food for two or more nodes. The sum of successes divided by the number of interactions results in an empirical estimate of $\rho\approx 0.82$.
@@ 6 Results
We have analysed the full fFPT distribution for homogenous populations of either scouts or recruits, using the same food placements (see Appendix 3 for specific sites) and the number of agents as in the empirical data (see Fig. 6). For all distributions, the mean and modal times were lower for the scouts than the recruits. With the exception of experiments two and six, the empirical times of first food item capture (black crosses) falls in a timescale which may be characterised by the mode of the theoretical distribution, showing that generally, the empirical time of food discovery corresponds to the direct regime of the theoretical distributions.
Figure 6: The first first-passage time distributions in Eq. (16), evaluated with the output of Eq. (22), for a population of scouts (red) and recruits (blue) modelling experiments one to nine in the honeycomb arena with hard reflecting walls. Physical target locations, $\bm{s}$, are given in Appendix 3, and the set of targets $\mathcal{T}$ is comprised of the three internal states within. Results for each experiment are shown in order (from one to nine) starting from the top left corner panel and ending at the bottom right corner panel, moving from one panel to the next one by rows. Crosses denote the empirical values for the food item found while red (scouts) and blue (recruits) dots indicate the mean value of the fFPT distributions. The $x$-axis has been scaled by a division of 60 to give results in minutes.
@@
With ant colonies facing scenarios with multiple food sources in nature, we have explored how switching behaviour, e.g., from scout to recruit, may affect search efficiency in a landscape with multiple targets (Fig. 7). For each experiment we assume a homogeneous population comprised of the number $\mathcal{N}$ ants observed experimentally in the laboratory set-up (written explicitly in Fig. 6), starting either as scouts or recruits. By assuming that the entire population may may modify searching statistics upon finding the first target or when returning back to the nest, we address the question of behavioural switching in two possible foraging strategies; (i) nest-to-patch foraging, i.e. strict central place foraging, in which ants go back to the nest before searching for the next target, and (ii) patch-to-patch foraging, where the ants use the previously discovered patch as a starting point to search for new food targets. The theoretical time for scenario (i) is defined as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{n}_{0};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})$, with $\bm{s}_{1}$ being the first found patch, and $\bm{s}_{2}$ the second one, while we define the theoretical time for scenario (ii) as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{s_{2}};\{1,\mathcal{N}\})$. Note that since the process is bounded by reflecting walls, the first first-passage probability is not symmetric with respect to the initial condition and target i.e., $F_{\bm{n}_{0}}(\bm{s},t;\{1,\mathcal{N}\})\neq F_{\bm{s}}(\bm{n}_{0},t;\{1,\mathcal{N}\})$.
@@
Figure 7: Relationship between the fFPT of the first patch (x-axis) and the second patch (y-axis). We considered two different foraging strategies: using previously found resources as starting points to search for new food targets (Patch to Patch, left panels) and strict central place foraging, wherein ants go back to the nest before searching for subsequent food targets (Nest to Patch, right panels). We allow the entire population to switch between scout (Sc.) and recruit (Re.) upon reaching the first target in the patch to patch case or when returning from the nest in the nest to patch scenario. As the theoretical values we show the mode of the fFPT, $\mathfrak{F}_{\bm{n}_{0}}(\bm{s};\{1,\mathcal{N}\})$ in red, and the first food capture time in blue. The numbers relate to the experimental setup, i.e. target placements and number of ants, given explicitly in Appendix 3 and Fig. (6), respectively. Due to the large emprical value in experiment two, we do not plot this case here.
@@
To interpret all experiments together, we show the overall trend in Fig. 7 where we see that scouts searching for the first patch reduced the time required for finding the targets, i.e. smaller modal fFPT, regardless of the foraging strategy (top panels), while switching from scout to recruit (Sc.-Re.) did not improve search efficiency in any of the setups. In agreement with previous results (Figs. 4 and 6), recruits starting the search resulted in larger modal fFPT irrespective of the context or the behavioural switching (Fig. 7, bottom panels). This was particularly evident in some spatial configurations, i.e. setups 3, 4 and 6, wherein the fFPT departed significantly from the empirical distribution.
@@ 7 Discussion
When collectively foraging, scouts and recruits are often identified in different ant species. They are commonly distinguished by their positions relative to the nest, but recent studies of the ant A. senilis have shown that scouts and recruits are not only located in different positions relative to the nest but also move pol2024liquid and behave martin2024 differently. The dynamics of these two types of ants in group or mass recruitment strategies are often complex and it is not easy to identify whether these are fixed or switching roles. In this chapter, by quantifying the microscopic movement parameters for each type, we studied the macroscopic consequences of locally different movement of both scouts and recruits in terms of their exploratory and search efficiency. In particular, we studied the mean number of sites visited and mean squared displacement for both ant roles, and also explored combinations of both types of ants in central place and non-central place foraging scenarios. The latter allowed us to assess differences in foraging efficiency and identify what are the more realistic scenarios when compared to empirical observations.
@@
The theoretical predictions in Fig. 7 offer indications that a central place foraging strategy is undertaken by the ants. These results suggest a temporal order in foraging, wherein ants first need to find a patch, convey the information to the rest of the colony on their way back to the nest to start patch exploitation, and then restart the search. This makes single-patch exploration-exploitation tradeoff be solved as a sequential process in time, where the exploration effort and the patch distance to the nest mainly drive the foraging efficiency. However, when there are multiple food patches, ant colonies may parallelise the exploitation as a function of the exploration effort of the colony, i.e. the number of scouts, and the initial positions from where scouting behaviour starts pol2024liquid .
@@
Moreover, Fig. 7 shows that scouts switching to recruit behaviour did not improve the colony performance when searching for multiple patches. In fact, in all exploration (Fig. 3) and search (Figs. 5 and 6) scenarios studied, the scouts always outperformed the recruits. This finding is supported by past theoretical investigations on square lattices showing that introducing persistence into the motion of a searcher can lower searching times in comparison to a diffusing motion tejedor2012optimizing . Since our results show that recruits are statistically close to being diffusive while scouts show persistent motion, one expects the scouts to be more efficient. Possible future work in this context includes using the honeycomb CRW results presented here to determine the optimum level of persistence on the honeycomb lattice to assess the overall efficiency of the scouts’ search.
@@
Since all our results point to recruits being less efficient explorers than scouts, we question whether recruits have any role in exploration. If diffusive or sub-diffusive searching by recruits around an area is important in detecting nearby food around already-discovered patches, recruits may have less of a role in explorative tasks. Because of their motion, recruits often end up aggregated, maintaining a minimum group cohesion in certain areas, which might be crucial for rapid information transfer and exploitation of known resources. To provide conclusive answers to these type of questions, further research is needed to explore the roles that movement statistics and switching behaviour play in exploration-exploitation tradeoffs in foraging.
@@
From a theoretical standpoint there has been a wealth of studies on the scaling of the MfFPT when the number of walkers is large and homogeneous weiss1983order ; lawley2020universal ; lawley2020probabilistic ; meerson2015mortality . Here, we have focused on experimentally relevant $\mathcal{N}$ and extended the theory to account for discrete time, the calculation of the splitting probability and also the possibility of a heterogeneous population comprised of walkers with different movement characteristics. As well as addressing the biologically motivated examples, we have utilised the theory to study some theoretical scenarios on the square lattice. We have uncovered (Fig. 4) oscillatory splitting probabilities when persistence is introduced into the population and also show that the number of persistent walkers in the population impacts the likelihood that each target in a multi-target search is found before the other. We have confirmed, in Fig. 4(b), that similarly to the one-walker case tejedor2012optimizing ; marris2024persistent setting the persistence length too high in a multi-walker search, may cause an increase in the MfFPT (Fig. 4(b)).
@@
To explain this aspect, consider that the MFPT may be thought of as a measure of the effectiveness of the direct trajectories godec2015optimization . With more effective walkers in the population, the chance of a direct trajectory being successful increases, which causes the MFPT to fall at a lower time as the mean becomes less skewed by long meandering trajectories. As a result, the fFPT distribution narrows and becomes more symmetric around the mode. If, on the other hand, a large subset of the population are ineffective, such as those that are highly persistent and do not explore the space efficiently, the tails of the distribution will be very long and the MfFPT will increase. In such extreme cases where the persistence length is much larger than the domain size we show that replacing diffusive walkers with persistent walkers in your population actually decreases search efficiency.
@@
We note that while the mean (and global mean) first-passage times condamin2005first ; condamin2007first ; benichou2014first ; tejedor2012optimizing can be useful in quantifying the overall efficiency of searches, more recent studies have suggested that the MFPT can often over-estimate the most-probable value for the first-passage time mattos2012first and often falls in the region associated with the indirect trajectories godec2015optimization . Here, the results of Fig. 6 show clearly that the experimental times in which the food is found are associated with the direct trajectories and are more reminiscent of the modal timescale than the mean of the fFPT distribution. This corroborates with what we see experimentally where trajectories in the exploration phase tend to be direct (Fig. 1). As exemplified by these trajectories, the ants perform foraging bouts, with looping trajectories of different lengths, which do not meander and explore the entire space. Instead, these trajectories extend from the nest outwards, and then, after an unsuccessful forage, return to the nest.
@@
This spatio-temporal patterns may explain why two of the nine theoretical predictions (exps. 2 and 6) in Fig. 6 underestimate the empirical findings and show the empirical time falling in a region of near zero probability of a first first-passage event. After an unsuccessful foraging bout the ant may return to, and remain in, the nest before performing another foraging excursion. Experimentally, the clock is not stopped if this occurs. This behavioural trait is not captured by the model, which, by construction, assumes that once a trajectory is started the walker will search indefinitely in the arena until they find the target site. Therefore, a natural extension to the model would be to include a mechanism for preferential returns to the initial condition, which may be included via a discrete space equivalent to the Smoluchowski-type diffusion equation kac1947random or via the inclusion of a process of resetting, whose discrete formulation has been recently derived das2022discrete .
@@
It is known that after the first food source is found, not only does interactions between ants increase, but the number of ants in the arena also increases cristin2024spatiotemporal . More specifically, it was observed that during the period between the first and second food source being found the number of ants in the system increased monotonically cristin2024spatiotemporal . Recognising a need to incorporate a varying number of ants into the fFPT calculations when studying foraging strategies, future work may also include the introduction of a dynamic population size into the fFPT, which is currently known only for continuous one-dimensional diffusion campos2024dynamic . Furthermore, the calculation of the fFPT to the second found food assumes the excursion from the nest to the first found patch is independent of the excursion from the first patch to the second (or back to the nest), an assumption which is clearly violated in practice. This highlights the need to develop a theory on sequential first-passage times, that is the first-passage probability to a target $\bm{s}_{2}$ conditioned on the trajectory first travelling through a previous target $\bm{s}_{1}$.

### v2

@@ 5.2 Experimentally relevant search on the Y-maze arena
We now consider a model which faithfully represents the domain size and boundaries; food and nest locations; and the number of ants in the system for each experimental realisation. Since the ant population in the arena was dynamic, with ants constantly entering and leaving the nest, we estimated the number of ants by considering how many ants were exploring the arena at the time when any food target was found (end of exploration phase) as the static number $\mathcal{N}$ for the model. The experimental set-up is modelled as a honeycomb lattice confined in a rectangular space, with hard, ‘bouncing’ reflecting walls, which causes persistent walkers to change direction at the boundary and continue its motion in the opposite direction marris2024persistent .
@@
To implement this zig-zag reflecting boundary condition via a fully analytical procedure is a formidable task when the movement is not diffusive. To proceed we use a semi-analytical iterative procedure gueneau2024siegmund ; marris2024persistent , which has been used to study persistent motion with complicated boundary conditions where the analytical representation of the occupation probability is not available. This procedure requires the decomposition of the movement dynamics into matrices that govern only the jumps from each state $i$ to every other state $j$, denoted by $\mathbb{M}(i,j)$. In our case this leads to 36 sparse transition matrices, of size $N_{1}N_{2}\times N_{1}N_{2}$. More concretely, the matrices that govern movement in the bulk have $1/N_{1}N_{2}$ ($\approx 0.31\%$) non-zero elements, while those governing movement on the boundary have $1/N_{1}^{2}N_{2}$ ($\approx 0.022\%$) or $1/N_{1}N_{2}^{2}$ ($\approx 0.0135\%$) non-zero elements. Hence, all matrices benefit from the use of sparse methods leading to efficient computation. See Appendix 3 for a detailed discussion of the transition matrices.
Encoded into $\mathbb{M}(i,j)$ is also the probability of absorption, i.e., the outgoing probabilities of the trap site, i.e., the site in which the first food source is found, sum up to $1-\rho$, where $\rho$ is the probability of absorption (finding the food) at the target. Experimentally, we estimated $\rho$ by measuring the result of each ant-food interaction. Namely, given an interaction event (i.e., an ant finds a food item), we consider a success (absorption) the ant collecting the food item, or a failure otherwise. Subsequent interactions were considered independent when the ant had walked away from the food for two or more nodes. The sum of successes divided by the number of interactions results in an empirical estimate of $\rho\approx 0.82$.
@@ 6 Results
We have analysed the full fFPT distribution for homogenous populations of either scouts or recruits, using the same food placements (see Appendix 4 for specific sites) and the number of agents as in the empirical data (see Fig. 6). For all distributions, the mean and modal times were lower for the scouts than the recruits. With the exception of experiments two and six, the empirical times of first food item capture (black crosses) falls in a timescale, which may be characterised by the mode of the theoretical distribution, showing that generally, the empirical time of food discovery corresponds to the direct regime of the theoretical distributions.
Figure 6: The first first-passage time distributions in Eq. (16), evaluated with the output of Eq. (22), for a population of scouts (red) and recruits (blue) modelling experiments one to nine in the honeycomb arena with hard reflecting walls. Physical target locations, $\bm{s}$, are given in Appendix 4, and the set of targets $\mathcal{T}$ is comprised of the three internal states within. Results for each experiment are shown in order (from one to nine) starting from the top left corner panel and ending at the bottom right corner panel, moving from one panel to the next one by rows. Crosses denote the empirical times for the food item found while red (scouts) and blue (recruits) dots indicate the mean value of the fFPT distributions. The $x$-axis has been scaled by a division of 60 to give results in minutes.
@@
With ant colonies facing scenarios with multiple food sources in nature, we have explored how switching behaviour, e.g., from scout to recruit, may affect search efficiency in a landscape with multiple targets (Fig. 7). For each experiment we assume a homogeneous population comprised of the number $\mathcal{N}$ ants observed experimentally in the laboratory set-up (written explicitly in Fig. 6), starting either as scouts or recruits. By assuming that the entire population may modify searching patterns upon finding the first target or when returning back to the nest, we address the question of behavioural switching in two possible foraging strategies; (i) nest-to-patch foraging, i.e., strict central place foraging, in which ants go back to the nest before searching for the next target, and (ii) patch-to-patch foraging, where the ants use the previously discovered patch as a starting point to search for new food targets. The theoretical time for scenario (i) is defined as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{n}_{0};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})$, with $\bm{s}_{1}$ being the first found patch, and $\bm{s}_{2}$ the second one, while we define the theoretical time for scenario (ii) as $\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{2}};\{1,\mathcal{N}\})=\mathfrak{F}_{\bm{n}_{0}}(\bm{s_{1}};\{1,\mathcal{N}\})+\mathfrak{F}_{\bm{s_{1}}}(\bm{s_{2}};\{1,\mathcal{N}\})$. Note that since the process is bounded by reflecting walls, the first first-passage probability is not symmetric with respect to the initial condition and target i.e., $F_{\bm{n}_{0}}(\bm{s},t;\{1,\mathcal{N}\})\neq F_{\bm{s}}(\bm{n}_{0},t;\{1,\mathcal{N}\})$.
@@
Figure 7: Relationship between the fFPT of the first patch (x-axis) and the second patch (y-axis). We considered two different foraging strategies: using previously found resources as starting points to search for new food targets (Patch to Patch, left panels) and strict central place foraging, wherein ants go back to the nest before searching for subsequent food targets (Nest to Patch, right panels). We allow the entire population to switch between scout (Sc.) and recruit (Re.) upon reaching the first target in the patch to patch case or when returning from the nest in the nest to patch scenario. As the theoretical values we show the mode of the fFPT, $\mathfrak{F}_{\bm{n}_{0}}(\bm{s};\{1,\mathcal{N}\})$ in red, and the first food capture time in blue. The numbers relate to the experimental setup, i.e., target placements and number of ants, given explicitly in Appendix 4 and Fig. (6), respectively. Due to the large emprical value in experiment two, we do not plot this case here.
@@
To interpret all experiments together, we show the overall trend in Fig. 7 where we see that scouts searching for the first patch reduced the time required for finding the targets, i.e., smaller modal fFPT, regardless of the foraging strategy (top panels), while switching from scout to recruit (Sc.-Re.) did not improve search efficiency in any of the setups. In agreement with previous results (Figs. 4 and 6), recruits starting the search resulted in larger modal fFPT irrespective of the context or the behavioural switching (Fig. 7, bottom panels). This was particularly evident in some spatial configurations, i.e., setups 3, 4 and 6, wherein the fFPT departed significantly from the empirical distribution.
@@ 7 Discussion
When collectively foraging, scouts and recruits are often identified in different ant species. They are commonly distinguished by their positions relative to the nest, but recent studies of the ant A. senilis have shown that scouts and recruits are not only located in different positions relative to the nest but also move pol2024liquid and behave martin2024 differently. The dynamics of these two types of ants in group or mass recruitment strategies are often complex and it is not easy to identify whether these are fixed or switching roles. In this chapter, we studied the macroscopic consequences of locally different movement of both scouts and recruits in terms of their exploratory and search efficiency. In particular, we studied the mean number of sites visited and mean squared displacement for both ant roles, and also explored combinations of both types of ants in central place and non-central place foraging scenarios. The latter allowed us to assess differences in foraging efficiency and identify what are the more realistic scenarios when compared to empirical observations.
@@
The theoretical predictions in Fig. 7 offer indications that a central place foraging strategy is undertaken by the ants. These results suggest a temporal order in foraging, wherein ants first need to find a patch, convey the information to the rest of the colony on their way back to the nest to start patch exploitation, and then restart the search. This makes single-patch exploration-exploitation tradeoff be solved as a sequential process in time, where the exploration effort and the patch distance to the nest mainly drive the foraging efficiency. However, when there are multiple food patches, ant colonies may parallelise the exploitation as a function of the exploration effort of the colony, i.e., the number of scouts, and the initial positions from where scouting behaviour starts pol2024liquid .
@@
Moreover, Fig. 7 shows that scouts switching to recruit behaviour did not improve the colony performance when searching for multiple patches. In fact, in all exploration (Fig. 3) and search (Figs. 5 and 6) scenarios studied, the scouts always outperformed the recruits. This finding is supported by past theoretical investigations on square lattices showing that introducing persistence into the motion of a searcher can lower searching times in comparison to a diffusing motion tejedor2012optimizing . Since our results show that recruits are statistically close to being diffusive while scouts show persistent motion, one expects the scouts to be more efficient. Possible future work in this context includes using the results on the honeycomb persistent random walk presented here to determine the optimum level of persistence on the honeycomb lattice to assess the overall efficiency of the scouts’ search.
@@
Since all our results point to recruits being less efficient explorers than scouts, we question whether recruits have any role in exploration. If diffusive or sub-diffusive searching by recruits around an area is important in detecting nearby food around already-discovered patches, recruits may have less of a role in explorative tasks. Because of their motion, recruits often end up aggregated, maintaining a minimum group cohesion in certain areas, which might be crucial for rapid information transfer and exploitation of known resources. To provide conclusive answers to these questions, further research is needed to explore the roles that different movement patterns and switching behaviours could play in exploration-exploitation tradeoffs while foraging.
@@
From a theoretical standpoint there has been a wealth of studies on the scaling of the MfFPT when the number of walkers is large and homogeneous lawley2020universal ; lawley2020probabilistic ; meerson2015mortality ; weiss1983order . Here, we have focused on experimentally relevant $\mathcal{N}$ and extended the theory to account for discrete time, the calculation of the splitting probability and also the possibility of a heterogeneous population comprised of walkers with different movement characteristics. As well as addressing the biologically motivated examples, we have utilised the theory to study some theoretical scenarios on the square lattice. We have uncovered (Fig. 4) oscillatory splitting probabilities when persistence is introduced into the population and also shown that the number of persistent walkers in the population impacts the likelihood that each target in a multi-target search is found before the other. We have confirmed, in Fig. 4(b), that similarly to the one-walker case marris2024persistent ; tejedor2012optimizing setting the persistence length too high in a multi-walker search, may cause an increase in the MfFPT (Fig. 4(b)).
@@
To explain this aspect, consider that the MFPT may be thought of as a measure of the effectiveness of the direct trajectories godec2015optimization . With more effective walkers in the population, the chance of a direct trajectory being successful increases, which causes the MFPT to fall at a lower time as the mean becomes less skewed by long meandering trajectories. As a result, the fFPT distribution narrows and becomes more symmetric around the mode. If, on the other hand, a large subset of the population are ineffective, such as those that are highly persistent and do not explore the space efficiently, the tails of the distribution will be very long and the MfFPT will increase. In such extreme cases where the persistence length is much larger than the domain size, we show that replacing diffusive walkers with persistent walkers in the population actually decreases search efficiency.
@@
We note that while the mean (and global mean) first-passage times benichou2014first ; condamin2005first ; condamin2007first ; tejedor2012optimizing can be useful in quantifying the overall efficiency of searches, more recent studies have suggested that the MFPT can often over-estimate the most-probable value for the first-passage time mattos2012first and often falls in the region associated with the indirect trajectories godec2015optimization . Here, the results of Fig. 6 show clearly that the experimental times in which the food is found are associated with the direct trajectories and are more reminiscent of the modal timescale than the mean of the fFPT distribution. This corroborates with what we see experimentally where trajectories in the exploration phase tend to be direct (Fig. 1). As exemplified by these trajectories, the ants perform foraging bouts, with looping trajectories of different lengths, which do not meander and explore the entire space. Instead, these trajectories extend from the nest outwards, and then, after an unsuccessful forage, return to the nest.
@@
These spatio-temporal patterns may explain why two of the nine theoretical predictions (exps. 2 and 6) in Fig. 6 underestimate the empirical findings and show the empirical time falling in a region of near zero probability of a first first-passage event. After an unsuccessful foraging bout the ant may return to, and remain in, the nest before performing another foraging excursion. Experimentally, the clock is not stopped if this occurs. This behavioural trait is not captured by the model, which, by construction, assumes that once a trajectory is started the walker will search indefinitely in the arena until they find the target site. Therefore, a natural extension to the model would be to include a mechanism for preferential returns to the initial condition, which may be included via a discrete space equivalent to the Smoluchowski-type diffusion equation kac1947random or via the inclusion of a process of resetting, whose discrete formulation has been recently derived das2022discrete ; barbinigiuggioli2024 .
@@
It is known that after the first food source is found, not only do interactions between ants increase, but the number of ants in the arena also increases cristin2024spatiotemporal . More specifically, it was observed that during the period between the first and second food source being found the number of ants in the system increased monotonically cristin2024spatiotemporal . Recognising a need to incorporate a varying number of ants into the fFPT calculations when studying foraging strategies, future work may also include the introduction of a dynamic population size into the fFPT, which is currently known only for continuous one-dimensional diffusion campos2024dynamic . Furthermore, the calculation of the fFPT to the second found food assumes the excursion from the nest to the first found patch is independent of the excursion from the first patch to the second (or back to the nest), an assumption which is clearly violated in practice. This highlights the need to develop a theory on sequential first-passage times, that is the first-passage probability to a target $\bm{s}_{2}$ conditioned on the trajectory first travelling through a previous target $\bm{s}_{1}$.